# Homophily-Driven Sanitation View for Robust Graph Contrastive Learning

### v2

@@ VII-A3 Baseline Models
To thoroughly validate the robustness of our proposed method, we have conducted a comparative analysis with several state-of-the-art graph contrastive learning models: BGRL [26], DGI [25], MVGRL [24], GRACE [3], GCA [4], ARIEL [5], and SPAN [6]. Among them, ARIEL and SPAN are known for their robustness in GCL. Meanwhile, it is worth noting that the defense methods proposed in [29, 30] require access to the clean graph, which is not available in a more practical setting such as ours, ARIEL and SPAN. To ensure a rigorous comparison, we have used the same hyperparameter settings as reported in the corresponding literature, where all the baselines are fine-tuned based on the performance of the validation set.
VII-B Robustness on Node Classification
@@ VII-B1 Against Mettack
Tab. III displays the accuracies of various GCL models for the semi-supervised node classification under different attacking powers. In particular, the attacking power is defined as the percentage $\frac{B}{|E|}$, where $B$ is the attack budget (number of modified edges) and $|E|$ is the total edge number of the clean graph. To cover most attacking scenarios, we have chosen the attacking power from $\{5\%,10\%,15\%,20\%\}$, which is consistent with the prior literature [12, 13].
@@
The results in Tab. III indicate that Mettack affects all GCL methods, although it is originally designed to attack vanilla GNNs. This is because the GCL models also rely on the GNN encoder to convert the graph data into low-dimensional embeddings, and thus, Mettack can indirectly compromise the quality of the GCL’s node embeddings.
@@
The second observation is that our proposed model outperforms the other baselines across various levels of attacking power. Moreover, the performance gap between our model and other baselines widens as the attacking power grows from $5\%$ to $20\%$. For example, on the Cora dataset, the performance gaps between our model with GRACE, ARIEL, GCA, SPAN, BGRL, DGI and MVGRL are $8.02\%$, $4.90\%$, $6.32\%$, $5.84\%$, $14.99\%$, $6.82\%$ and $3.87\%$ when the attacking power is $5\%$. The corresponding margins under $20\%$ attacking power are $40.17\%$, $36.47\%$, $33.11\%$, $33.57\%$, $35.79\%$, $25.26\%$, $16.74\%$. This can be attributed to the fact that as the attacking power grows, more malicious heterophilous links are inserted. However, the sanitizer in our proposed model can effectively eliminate a greater number of malicious edges from the heavily poisoned graph, thereby improving the quality of node embeddings and making them more distinguishable across different classes.
Thirdly, it is observed that in certain scenarios, our model cannot surpass other strong baselines on clean graphs. For instance, on the CiteSeer dataset, the node classification accuracy of our model is slightly lower than that of GCA and MVGRL by $0.57\%$ and $0.55\%$. This phenomenon is common in the field of adversarial machine learning and can be attributed to the trade-off between expressive power and the robustness of the model [45, 46].
VII-B2 Against CLGA
Tab. IV presents the robustness results under the unsupervised attack CLGA. Similar to its performance under Mettack, our model outperforms other baselines across various levels of attacking power. However, since CLGA is an unsupervised attack, the degree of decreasing node classification accuracy for GCL models is not as significant as in Mettack, which specifically targets the test set.
@@ VII-B2 Against CLGA
We note that MVGRL performs well against CLGA and even outperforms our method when the attacking power is equal to $5\%$ on the Cora dataset. This is because MVGRL does not use the infoNCE loss for training, which is the attack objective of CLGA. Nevertheless, our model outperforms MVGRL on the Cora dataset when the attacking power is larger or equal to $10\%$. Overall, the proposed homophily-driven learnable sanitation view effectively promotes the robustness of the vanilla GCL framework by automatically detecting and pruning potential heterophilous edges in the poisoned graph, as demonstrated by the robust performance results shown in Tab. III and Tab. IV.
VII-C Robustness on Graph Clustering
In this section, we examine the adversarial robustness of GCL models on the graph clustering downstream task, with a focus on the unsupervised graph structural attack method CLGA. This attack method is particularly relevant since it can significantly degrade the quality of node embeddings, irrespective of the downstream task. The attacking scenario is the same as described in Section VII-B, and the experiment results are presented in Fig. 4.
Firstly, we investigate the attack performance of CLGA against the graph clustering task by observing the degeneration percentage of GCA’s clustering performance (CLGA chooses GCA as its representative surrogate model). It is observed that for the four graph data listed in Fig. 4, the NMI score of the GCA model indeed decreases as the graph attacker increases the attacking power. For example, compared to the clean graph, the decreasing percentages on the poisoned Cora dataset are $22.22\%$, $24.07\%$, $37.04\%$, $40.74\%$ when the attacking power increases from $5\%$ to $20\%$. This phenomenon demonstrates that CLGA can successfully attack the graph clustering task by indirectly attacking the infoNCE loss of the graph contrastive learning framework.
The second key finding is that our proposed model achieves the best performance in most cases with various attacking powers ranging from $5\%$ to $20\%$. For instance, under the attacking power equal to $10\%$, our method outperforms GCA and SPAN around $19.35\%$ and $25.81\%$ for CiteSeer. Similar to the node classification task, our method cannot achieve the best clustering performance on the clean graph. This bottleneck is still due to the trade-off between the expressive power and the robust model. Alternatively, SPAN can perform well in certain scenarios and even outperform our method on the Photo data when the attacking power is equal to $15\%$ demonstrating that SPAN has intrinsic robustness property on the graph clustering task. The reason is that the spectrum-invariance-based augmentation of SPAN can inherently preserve the clustering consistency to some extent. Specifically, the graph spectrum is a summary of graph structural properties such as clustering [47, 48]. SPAN enhances the robustness of graph clustering by selecting two augmentation views to simultaneously maximize and minimize the spectrum distance of the original graph and corrupted graph in the augmentation view while maintaining consistency on the embedding space between the two augmentation views from the contrastive loss. However, the fact that our method performs better than SPAN in most cases demonstrates that the homophily-driven sanitation view of sanitizing the potential inter-class links is a more appropriate defense method to promote the robustness of the graph clustering task.
(a) CLGA
(b) Mettack
@@ VII-C Robustness on Graph Clustering
(a) CLGA
(b) Mettack
Figure 6: Sensitivity analysis on $\eta$.
(a) GCHS-CLGA
(b) GCHS-Mettack
(c) SPAN-CLGA
(d) SPAN-Mettack
(e) ARIEL-CLGA
(f) ARIEL-Mettack
Figure 7: Scatterplots for GCHS, SPAN and ARIEL.
(a) CLGA
(b) Mettack
Figure 8: Performances on GRV.
VII-D Ablation Study
In Sec. VI-B3, we concluded that both infoNCE loss and the graph homophily $\delta_{x}$ can serve as appropriate supervisors and lead to high-quality sanitized node embeddings. In this section, we present the ablation study to verify the effectiveness of these two vital components. Particularly, we introduce two model variants as follows:
@@ VII-D Ablation Study
  • GCHS w.o. $\mathcal{L}_{info}$: we remove the infoNCE loss in Eqn. 16 and first train the sanitation view to optimize $\delta_{x}$. Then, we train the GCL model with the pre-trained sanitation view.
@@
  • GCHS w.o. $\delta_{x}$: we remove the $\delta_{x}$ term in Eqn. 16 and only optimize the infoNCE loss in GCHS.
@@
The results are presented in Fig. 5. Firstly, comparing the robust performance of GCHS w.o. $\mathcal{L}_{info}$, GCHS w.o. $\delta_{x}$ and GRACE, we find out that both optimizing the infoNCE loss and the graph homophily $\delta_{x}$ can enhance the robustness of the GCL model since in most cases the two model variants outperform GRACE. Secondly, the comparison of GCHS w.o. $\mathcal{L}_{info}$ and GCHS w.o. $\delta_{x}$ under Mettack reveals that optimizing $\delta_{x}$ provides more powerful robustness than optimizing $\mathcal{L}_{info}$. The reason is that Mettack will maliciously inject more inter-class links than CLGA and directly optimizing the graph homophily will have a more significant sanitation gain than indirectly optimizing the task-related loss function. Thirdly, the observation that GCHS outperforms both model variants demonstrate that considering both components is necessary to achieve the best sanitized node embeddings.
VII-E Sensitivity Analysis
@@
We examine the sensitivity of GCHS to the hyperparameter $\eta$ in Eqn. 16. To be specific, we report the best pseudo normalized cut loss $\mathcal{L}_{pnc}^{t^{*}}$ and the node classification accuracy with attack power equal to $10\%$ for CLGA and Mettack as exemplar and present the results in Fig. 6. Moreover, we tune the hyperparameter $\eta$ ranged in $[0,0.0001,0.001,0.01,0.1,1,3,5].$ The experiments demonstrate that too large and too small $\eta$ may lead to sub-optimal performance on GCHS under different attacking scenarios. Alternatively, for all the cases listed in Fig. 6 $\eta=1$ can achieve the minimum $\mathcal{L}_{pnc}^{t^{*}}$ (coincide with the best node classification accuracy).
VII-F Embeddings Visualization
@@
We provide the node embeddings visualization for GCHS, SPAN and ARIEL (robust GCLs) to qualitatively evaluate the quality of node embeddings under different attacking scenarios (CLGA and Mettack with attacking power equal to $20\%$) in Fig. 7. The visualization is depicted on the transformed node embeddings generated by GCHS, SPAN and ARIEL via t-SNE [49]. Clearly, the node embeddings for GCHS are more discriminative than SPAN and ARIEL and the clusters of GCHS are more cohesive than the others.
VII-G Robustness on GRV
@@
To illustrate the robustness of GCHS from an information-theoretic perspective, we utilize GRV [29, 30] mentioned in Def. 1 to indirectly quantify the GCL’s robustness. It is worth noting that GRV measures the difference between the mutual information of the node embeddings for the clean graph and the poisoned graph. Here we utilize the infoNCE loss to approximate the mutual information. We emphasize that we only use the clean graph’s information on computing GRV instead of training the robust model and hence prevent the information leakage. We compare GCHS with SPAN and ARIEL in this part and report the GRVs for each robust model in Fig. 8. A lower GRV indicates better robustness. It is observed that GCHS achieves the minimum GRV for both CLGA and Mettack, demonstrating that the node embeddings of GCHS are more robust than those of the other models from an information-theoretic perspective.
VII-H Time Cost
Table V: Time cost (s).
Dataset BGRL DGI MVGRL GRACE GCA ARIEL SPAN GCHS
@@ VII-H Time Cost
Tab. V presents the time required for each iteration of the training phase of the GCL models. The last three columns show the time cost of using the learnable augmentation views. It is evident that the GCL model with learnable augmentation views requires more computation time than the others. In addition to that, our model consumes the least amount of computation time compared to the other two models with learnable augmentation views.
VIII Conclusion
In this paper, we investigate the adversarial robustness of the GCL model in the face of graph structural attacks. This research proposes a novel homophily-driven, learnable sanitation view that enhances the GCL model’s robustness by automatically eliminating potential malicious edges in the augmentation view. To achieve this, we choose the objective as the weighted sum of the infoNCE loss with the graph homophily $\delta_{x}$ to jointly optimize the parameters in the sanitation view and GNN encoder. The Gumbel-Softmax re-parametrization technique is employed to overcome the non-differentiable mapping issue during the sampling phase. Besides, we novelly craft a pseudo normalized cut loss to easily determine the best choice of the hyperparameter $\eta$ without avoiding the unsupervised setting of the GCL framework during training. Extensive experiments demonstrate that our method can prominently boost the robustness of the GCL model.

### v3

@@ VI-B Baselines
To thoroughly validate the robustness of our proposed method, we have conducted a comparative analysis with several state-of-the-art graph contrastive learning models: BGRL [45], DGI [41], MVGRL [40], GRACE [3], GCA [4], ARIEL [5], and SPAN [6]. Among them, ARIEL and SPAN are known for their robustness in GCL. Meanwhile, it is worth noting that the defense methods proposed in [35, 58] require access to the clean graph, which is not available in a more practical setting such as ours, ARIEL and SPAN. We list the brief introduction of these baselines as follows:
  • BGRL [45] It introduces the Bootstrapped Graph Latents technique to improve the scalability of graph contrastive learning effectively. It contrasts the two embeddings from distinct encoders rather than augmentation views, in which one encoder is the exponential moving average of the other.
  • DGI [41] It relies on maximizing the mutual information between patch representations and corresponding high-level summaries of graphs, such as the readout mapping of the node embeddings.
  • MVGRL [40] It learns node and graph level representations by contrasting low-dimensional embeddings from two augmented views of graphs, including first-order neighbors and a graph diffusion.
  • GRACE [3] It is an unsupervised graph representation learning that generates two augmentation views by corruption and learns node representations by maximizing the agreement of node representations in these two views.
  • ARIEL [5] it leverages the adversarial training strategy to construct the adversarial view with information regularization to achieve adversarial robustness.
  • GCA [4] It is a graph contrastive learning framework that employs more complex adaptive augmentation schemes such as degree centrality, eigenvector centrality, and PageRank centrality.
  • SPAN [6] It introduces two augmentation views to minimize and maximize the spectrum of the augmented graphs to preserve the spectrum invariance of the GCL.
  • PiGCL [47] It is a graph contrastive learning framework that dynamically captures the implicit conflicts from the negative pairs during training by detecting the gradient of representation similarities to eliminate the conflict phenomenon caused by the InfoNCE loss.
  • SPAGCL [46] It is a graph contrastive learning framework that combines the node similarity-preserving view and an adversarial view to improve the node similarity during the adversarial training.
VI-C Robustness on Node Classification
@@ VI-C1 Against Mettack
Tab. II displays the accuracies of various GCL models for the semi-supervised node classification under different attacking powers. In particular, the attacking power is defined as the percentage $\frac{B}{|E|}$, where $B$ is the attack budget (number of modified edges) and $|E|$ is the total edge number of the clean graph. To cover most attacking scenarios, we have chosen the attacking power from $\{5\%,10\%,15\%,20\%\}$, which is consistent with the prior literature [18, 19].
@@
The results in Tab. II indicate that Mettack affects all GCL methods, although it was originally designed to attack vanilla GNNs. This is because the GCL models also rely on the GNN encoder to convert the graph data into low-dimensional embeddings, and thus, Mettack can indirectly compromise the quality of the GCL’s node embeddings.
@@
The second observation is that our proposed model outperforms the other baselines across various levels of attacking power. Moreover, the performance gap between our model and other baselines widens as the attacking power grows from $5\%$ to $20\%$. For example, on the Cora dataset, the performance gaps between our model with GRACE [3], ARIEL [5], GCA [4], SPAN [6], BGRL [45], DGI [41] and MVGRL [40] are $8.02\%$, $4.90\%$, $6.32\%$, $5.84\%$, $14.99\%$, $6.82\%$ and $3.87\%$ when the attacking power is $5\%$. The corresponding margins under $20\%$ attacking power are $40.17\%$, $6.47\%, 33$$11\%, 33.$$3.57\%$, $35.79\%$, $25.26\%$, $16.74\%$. This can be attributed to the fact that as the attacking power grows, more malicious links are inserted. However, the sanitation view in our proposed model can effectively eliminate a greater number of malicious edges from the heavily poisoned graph, thereby improving the quality of node embeddings and making them more distinguishable across different classes.
VI-C2 Against CLGA
@@
Tab. III presents the robustness results under the unsupervised attack CLGA. Similar to its performance under Mettack, our model outperforms other baselines across various levels of attacking power. However, since CLGA is an unsupervised attack, the degree of decreasing node classification accuracy for GCL models is not as significant as in Mettack, which specifically targets the test set. We note that MVGRL performs well against CLGA and even outperforms our method when the attacking power is equal to $5\%$ on the Cora dataset. This is because MVGRL does not use the InfoNCE loss for training, which is the attack objective of CLGA. Nevertheless, our model outperforms MVGRL on the Cora dataset when the attacking power is larger than or equal to $10\%$. Overall, the proposed learnable sanitation view effectively promotes the robustness of the vanilla GCL framework by automatically detecting and pruning potential malicious edges in the poisoned graph, as demonstrated by the robust performance results shown in Tab. II and Tab. III.
((a))
((b))
@@
((a))
((b))
Figure 8: Evaluation of the adversarial robustness of robust GCLs based on GRV for different attacking scenarios.
VI-D Ablation Study
In this section, we present the ablation study to verify the effectiveness of two vital components of the loss function defined in Eqn. 13, corresponding to restoring the mutual information between the graph and its representations and the graph homophily respectively. In Particular, we introduce two model variants as follows:
@@ VI-D Ablation Study
  • GCIR w.o. $\mathcal{L}_{info}$: we remove the InfoNCE loss in Eqn. 13 and first train the sanitation view to optimize $\delta_{x}$. Then, we train the GCL model with the pre-trained sanitation view.
@@
  • GCIR w.o. $\delta_{x}$: we remove the $\delta_{x}$ term in Eqn. 13 and only optimize the InfoNCE loss in GCIR.
@@
The results are presented in Fig. 7. Firstly, comparing the robust performance of GCIR w.o. $\mathcal{L}_{info}$, GCIR w.o. $\delta_{x}$ and GRACE, we find out that both optimizing the InfoNCE loss and the graph homophily $\delta_{x}$ can enhance the robustness of the GCL model since in most cases the two model variants outperform GRACE. Secondly, the comparison of GCIR w.o. $\mathcal{L}_{info}$ and GCIR w.o. $\delta_{x}$ under CLGA reveals that restoring the diminished mutual information will be more effective than the graph homophily. Thirdly, the comparison of GCIR w.o. $\mathcal{L}_{info}$ and GCIR w.o. $\delta_{x}$ under Mettack indicates that restoring the graph homophily will be more effective than the mutual information. The reason is probably that Mettack is such attacks that degenerate the GNN’s performance, which does not directly rely on information theory.
VI-E Sensitivity Analysis
@@
In this section, we discuss the tuning strategy of the vital hyperparameter $\eta$ in Eqn. 13. To mimic the real environment, it is assumed that the defender cannot acquire the attacking scenarios (such as the attack type, attack degree, and victim nodes or edges and etc.), and the label information of the given graph. That is, the defender can only utilize the attribute matrix, the suspected perturbed adjacency matrix, and the node embeddings of GCL to determine the best choice of $\eta$. To tackle this issue, as previously mentioned in Sec. V-C, we craft a pseudo normalized cut loss $\mathcal{L}_{pnc}^{*}$ to tune the vital hyperparameter $\eta$. To be specific, we report the best pseudo normalized cut loss $\mathcal{L}_{pnc}^{t^{*}}$ and the node classification accuracy with attack power equal to $10\%$ for CLGA and Mettack as exemplar and present the results in Fig. 6. Moreover, we tune the hyperparameter $\eta$ ranged in $[0,0.0001,0.001,0.01,0.1,1,3,5].$ The experiments demonstrate that too large and too small $\eta$ may lead to sub-optimal performance on GCIR under different attacking scenarios. Alternatively, for all the cases listed in Fig. 6 $\eta=1$ can achieve the minimum $\mathcal{L}_{pnc}^{t^{*}}$ (coincide with the best node classification accuracy).
((a))
((b))
((c))
((d))
((e))
((f))
Figure 9: Scatterplots for GCIR, SPAN and ARIEL.
VI-F Node Embeddings Visualization
@@
We provide the node embeddings visualization for GCIR, SPAN, and ARIEL (robust GCLs) to qualitatively evaluate the quality of node embeddings under different attacking scenarios (CLGA and Mettack with attacking power equal to $20\%$) in Fig. 9. The visualization is depicted on the transformed node embeddings generated by GCIR, SPAN, and ARIEL via t-SNE [59]. Clearly, the node embeddings for GCIR are more discriminative than SPAN and ARIEL and the clusters of GCIR are more cohesive than the others.
VI-G Robustness on Graph Representation Vulnerability
In this section, we introduce the concept of graph representation vulnerability and its relationship with the evaluation of the GCL’s robustness. We then test the robustness of the robust GCLs based on the defined metric from an information-theoretical perspective.
Definition 5 (GRV [35, 58]).
Definition 6 (Defense of GCL).
where $I(\cdot,\cdot)$ represents the mutual information, $f_{\theta}(\cdot)$ represents the GNN encoder parametrized by weights $\theta$.
@@ VI-G Robustness on Graph Representation Vulnerability
We then validate the robust performances of our proposed method with two robust baselines from an information-theoretical perspective. Specifically, we utilize GRV [35, 58] mentioned in Def. 5 to indirectly quantify the GCL’s robustness. It is worth noting that GRV measures the difference between the mutual information of the node embeddings for the clean graph and the poisoned graph. Here we utilize the InfoNCE loss to approximate the mutual information. We emphasize that we only use the clean graph’s information on computing GRV instead of training the robust model and hence prevent the information leakage. We compare GCIR with SPAN and ARIEL in this part and report the GRVs for each robust model in Fig. 8. A lower GRV indicates better robustness. It is observed that the GCIR achieves the minimum GRV for both CLGA and Mettack, demonstrating that the node embeddings of GCIR are more robust than those of the other models from an information-theoretical perspective.
VI-H Time Cost
TABLE IV: Training time per epoch (s).
Dataset BGRL DGI MVGRL GRACE GCA ARIEL SPAN GCIR
@@ VI-H Time Cost
Tab. IV presents the time required for each iteration of the training phase of the GCL models. The last three columns show the time cost of using the learnable augmentation views. It is observed that the GCL model with learnable augmentation views requires more computation time than the others. Moreover, our model consumes the least amount of computation time compared to the other two models with learnable augmentation views.
VI-I Discussions on Heterophilic Graphs
((a))
((b))
Figure 10: Models’ Robustness over Heterophilic Graphs.
In this section, we investigate the adversarial robustness of graph contrastive learning methods on heterophilic graphs. We take two typical heterophilic graphs, i.e., Chameleon and Squirrel [60] as examplars. Similarly, we deploy Mettack to inject adversarial noises into the clean heterophilic graphs with different attacking powers and evaluate the adversarial robustness of GCLs on the poisoned graphs. The experimental results in Fig. 10 demonstrate that our proposed method can still consistently outperform other baselines on heterophilic graphs. It is worth noting that the hyperparameter $\eta$ in Eqn. 13 is tuned to be zero under this case. It is due to the fact that the feature smoothness is invalid for heterophilic graphs whose node pairs are dissimilar in the majority. Hence, the proposed graph contrastive learning framework solely relies on the information restoration mechanism to achieve robustness. In future work, we will follow this research line and investigate how to further enhance the adversarial robustness of the graph contrastive learning framework over heterophilic graphs while maintaining its robustness over homophilic graphs.
VII Conclusion
In this paper, we investigate the adversarial robustness of the GCL against the graph structural attacks from an information-theoretical perspective. The vulnerability analysis offers the clue that apart from the conventional observation that structural attacks tend to diminish the graph homophily, these attacks also degenerate the mutual information estimation between the graph and its representations, which are the cornerstone of the powerful representation learning of GCL. Based on our findings, we propose GCIR with a learnable sanitation view to restore the diminished mutual information and graph homophily after attacks and thus achieve adversarial robustness. The whole framework is trained in an end-to-end learning manner. Besides, we novelly craft an unsupervised tuning strategy to easily determine the best choice of the GCL’s vital hyperparameter without avoiding the fully unsupervised setting during the training of GCL. Extensive experiments demonstrate that our method can prominently boost the robustness of the GCL model.
Acknowledgement
This research was partly supported by the Hong Kong RGC Project (No. PolyU25210821).